# GeoSDF: Plane Geometry Diagram Synthesis via Signed Distance Field

### v1

@@ Abstract.
Plane Geometry Diagram Synthesis has been a crucial task in computer graphics, with applications ranging from educational tools to AI-driven mathematical reasoning. Traditionally, we rely on computer tools (e.g., Matplotlib and GeoGebra) to manually generate precise diagrams, but it usually requires huge, complicated calculations cost. Recently, researchers start to work on learning-based methods (e.g., Stable Diffusion and GPT4) to automatically generate diagrams, saving operational cost but usually suffering from limited realism and insufficient accuracy. In this paper, we propose a novel framework GeoSDF to automatically generate diagrams efficiently and accurately with Signed Distance Field (SDF). Specifically, we first represent geometric elements (e.g., points, lines, circles) in the SDF, then construct a series of constraint functions to represent geometric relationship, next we optimize such constraint functions to get an optimized field of both elements and constrains, finally by rendering the optimized field we can obtain the synthesized diagram. In our GeoSDF, we define a symbolic language to easily represent geometric elements and those constraints, and our synthesized geometry diagrams can be self-verified in the SDF, ensuring both mathematical accuracy and visual plausibility. In experiments, our GeoSDF synthesized both normal high-school level and IMO-level geometry diagrams. Through both qualitative and quantitative analysis, we can see that synthesized diagrams are realistic and accurate, and our synthesizing process is simple and efficient. Furthermore, we obtain a very high accuracy of solving geometry problems (over 95% while the current SOTA accuracy is around 75%) by leveraging our self-verification property. All of these demonstrate the advantage of GeoSDF, paving the way for more sophisticated, accurate, and flexible generation of geometric diagrams for a wide array of applications. Based on the GeoSDF, we have developed a tool for plan geometry diagram synthesis and will provide the link after the paper review.
Plane Geometry Diagram Synthesis, Plane Geometry Problems, Math Reasoning, Signed Distance Field
submissionid: 1816journal: TOGccs: Theory of computation Computational geometryccs: Computing methodologies Representation of mathematical objectsccs: Computing methodologies Optimization algorithms
1. Introduction
Figure 1. Comparison of Geometric Diagram Synthesis Methods. This figure illustrates the distinct approaches to generating plane geometric diagrams. Manual systems require tedious coordinate and relation input (left side for each diagram), involving extensive numerical calculations. Rule-based systems (e.g., R-COT (Kazemi et al., 2024)), while structured, struggle with novel, complex, or highly precise mathematical relationships due to their reliance on predefined templates. Model-based systems DALL-E 3 (OpenAI, 2024) and GPT4o lack the precision and accurate mathematical synthesis. In contrast, our method synthesizes novel and complex diagrams accurately from elements and constraints without manual calculation. All diagrams are synthesised from the IMO problem set.
@@ 1. Introduction
The study of geometry problems is fundamental across numerous disciplines, serving as a cornerstone for advancements in computer graphics, computational geometry, robotics, architectural design, and even theoretical mathematics. The emergence of large language models (LLMs) has introduced new approaches for addressing Plane Geometry Problems (PGPs) (Shi et al., 2024) by understanding both textual problem descriptions and their corresponding geometric diagrams (Gao et al., 2025; Cai et al., 2024; Xia et al., 2025; Xu et al., 2024a; Lu et al., 2024; Zhang et al., 2024a, b). Despite significant efforts to enhance the reasoning capabilities of LLMs, the performance of models remains limited by the scarcity of high-quality datasets containing annotated plane geometry diagrams (Zhang et al., 2023b). Improvements in text-based inputs alone, like G-LLaVA (Gao et al., 2025), are insufficient to compensate for the absence of rich, well-structured graphical representations (Xu et al., 2024b; Ning et al., 2025). As a result, suboptimal performance appears in model-driven solutions (Deng et al., 2024). Therefore, our work aims to design an approach to synthesize high-quality plane geometry diagrams.
@@
In this paper, we propose a novel framework to automatically synthesize precise Geometry diagrams by leveraging Signed Distance Field (SDF), termed as GeoSDF. In the theory of SDF, it defines a scalar field that assigns to each point in space a signed distance to the nearest surface of a geometric object (Yariv et al., 2020), thus being differentiable to be optimization. In our proposed GeoSDF, the textual problem description is first parsed into a set of mathematically well-defined components, which we refer to as elements. These elements are then governed by a set of geometric constraints derived from the problem’s underlying mathematical relationships, such as perpendicularity, parallelism, or incidence. Unlike previous methods, our GeoSDF allows these relationships to be quantitatively expressed—for instance, through the distance from a point to a line—enabling precise and verifiable encoding of geometric structure. These constraints are then incorporated into a loss function, effectively translating the mathematical conditions into an optimization objective. By leveraging gradient descent, GeoSDF iteratively refines the positions of elements in the diagram to satisfy complex constraints, allowing for the generation of accurate and mathematically valid geometric figures. This formulation supports both quantifiable relations and the derivation of exact solutions, addressing key limitations of existing diagram generation methods.
In experiments, we synthesize both normal high-school level and IMO-level geometry diagrams, and both qualitative and quantitative analyses demonstrate that the synthesized geometry diagrams are realistic and accurate, verifying the effectiveness. Furthermore, by leveraging the self-verification property, GeoSDF can also be directly applied in solving geometry problems and achieve a very high accuracy, significantly outperforming previous learning based methods, including advanced LLMs such as GPT and Gemini. For example, we can achieve an accuracy of 95.9% on GeoQA dataset (Chen et al., 2021), which is obviously higher than the current state-of-the-art accuracy (only around 75%). To be noted, our GeoSDF is easy to implement, and the synthesis is efficient. Based on the GeoSDF, we have developed a tool for plan geometry diagram synthesis. The contributions of this paper can be summarized as follows
@@
  • We introduce GeoSDF, a novel framework that synthesizes plane geometry diagrams accurately. It integrates both structural elements by SDF representation and mathematical constraints within an optimization framework.
  • As a key feature of GeoSDF, all geometric relations are inherently quantifiable. The constraint satisfaction and diagram quality can be explicitly evaluated.
  • The utilization of quantifiable geometric relations enables our GeoSDF can directly infer accurate solutions to solve geometry problems.
  • Extensive experiments demonstrate that diagrams synthesized by GeoSDF are highly accurate and enhance the capabilities of AI models in geometric reasoning.
2. Related Works
2.1. Plane Geometry Problem Solving
@@
Early approaches to plane geometry problem (PGP) solving relied on manually defined rules applied to small datasets (Seo et al., 2015, 2014), resulting in poor generalization. Neural network-based models, such as NGS (Chen et al., 2021) and DPE-NGS (Chen et al., 2021), introduced visual question answering and specialized program generation, but still struggle with coarse-grained diagram understanding (Ning et al., 2023). Symbolic reasoning methods, including Inter-GPS (Lu et al., 2021) and FormalGeo (Zhang et al., 2023b), use complex rule-based systems to interpret formal languages parsed from diagrams and text. However, their performance is constrained by limited datasets and parameter sizes, and they do not produce natural language solutions. Recent advances in multi-modal large language models (MLLMs), such as G-LLaVA (Gao et al., 2025) and Geo-LLaVA (Xu et al., 2024b), fine-tune based models (Liu et al., 2023) to generate natural language solutions. These approaches benefit from data augmentation techniques (e.g., Geo170K via GPT-3.5 (Gao et al., 2025)), but often prioritize textual information over the diversity and complexity of geometry diagrams (Zhuang et al., 2024). While methods like GeoX (Xia et al., 2025) and R-CoT (Deng et al., 2024) have made progress in diagram understanding, generating accurate, diverse, and controllable geometric diagrams remains a fundamental challenge for advancing PGP solving.
2.2. Plane Geometry Diagram Synthesis
@@
Plane geometry diagram synthesis is critical for evaluating problem-solving systems but is underexplored due to complexity (Rombach et al., 2022). Current rule-based methods, including GeomVerse (Kazemi et al., 2024) and MAVIS (Zhang et al., 2024c), stitch predefined shapes along edges. This limits them, however, from generating diagrams with inscribed relationships (e.g., a triangle in a circle), impacting realism. Despite R-CoT’s (Deng et al., 2024) advancements with inscribed elements, existing approaches remain constrained by predefined shapes and rules, preventing synthesis from textual requirements. Our proposed GeoSDF overcomes this by synthesizing flexible plane geometry diagrams directly from given text constraints.
3. Methodology
In this section, we introduce the method for utilizing geometric elements and constraints to generate diagrams that fulfill all specified constraints. The SDF and geometric terms will be introduced in Sec. 3.1 while Section 3.2 describes our method in details consisting of four steps.
Figure 2. The visualization of the optimization process. We first randomly initialize the geometry elements and then optimize the SDF by geometry constraints to the synthesized geometry diagram. The upper two examples are selected from FormalGeo-IMO (Zhang et al., 2023b), and the lower two examples are from IMO 1959 Question 5 and IMO 2021 Question 4.
3.1. Preliminary
Signed Distance Field. A field in mathematics is a function that assigns a value (scalar or vector) to each point in a space. The space provides the coordinate system and the notion of distance necessary. The Signed Distance Field assigns a signed distance to the nearest point on the boundary of the shape for every point in this space. In mathematics, the SDF $F$ is a scalar field that assigns a value $s$ to every position $x$ in the space by
(1) $F(x)=s,\quad x\in\mathbb{R}^{2},\,s\in\mathbb{R},$
The value $s$ indicates the distance from the position $x$ to the shape $\Omega\in\mathbb{R}^{2}$ in the two-dimensional plane space. The distance between a point $x$ and the $\partial\Omega$ boundary of the shape is hereby defined as
$d(x,\partial\Omega)=\inf\limits_{y\in\partial\Omega}d(x,y),$
Elements and Constraints in geometry diagrams. In geometry, the fundamental elements in 2D Euclidean geometry are Points, Lines, Circles, and Planes. Constraints are conditions or relationships imposed on geometric elements. These constraints restrict the possible configurations or properties of the elements, leading to specific geometric diagrams or relationships. Constraints can be broadly classified into four categories (Berg et al., 2008): 1) Incidence Constraints: These specify that one geometric element lies on another. 2) Metric Constraints: These involve measurements such as distances, angles, and areas. 3) Relative Position and Orientation Constraints: These define how geometric elements are positioned and oriented with respect to each other. 4) Topological Constraints: These describe the connectivity and arrangement of geometric elements, focusing on properties that are preserved under continuous deformations.
3.2. Plane Geometry Diagram Synthesis
@@
For synthesizing plane geometry diagrams, our GeoSDF framework mainly consists of four steps as shown in Fig.3.
@@
Figure 3. The GeoSDF Pipeline with an example IMO diagram synthesis process. (a) The description of the figure in natural language. (b) The representation of geometric elements by our symbolic language. (c) The visualization of initialized elements field. (d) Constraints of elements in our SDF. (e) Visualization of the optimized fields. (f) The rendering diagrams. Some parts are omitted due to space limitations.
(1) Representation of Geometric Elements. We use the functions to represent basic geometric elements $E$ in the Signed Distance Field. We define the four types of elements: point, segment, line, and circle. The function defines the distance from each point in space to the shape. By compositing the basic elements, we could synthesize the complex shapes in Fig. 3 (b). These functions are parameterized by continuous variables (e.g., coordinates of points, radii of circles, or angles between lines) to control the attributes. In the initial state, we construct the parameters randomly. We sample the SDF by pixel grid directly at the initial state, as shown the Fig. 3 (c).
(2) Constraints of the Elements. Once the basic elements are represented, we impose geometric constraints, which define the relationships between these elements, as optimization targets. We define the constraint functions by using a symbolic language, shown in Tab. 1. These constraints are a set of differentiable functions, $C=\{c_{i}(E)=0\}_{i=1}^{m}$, where each $c_{i}(E)$ represents a specific constraint with elements set $E=\{e_{1},e_{2},\dots,e_{N}\}$ and $m$ constraints in total. The goal is to find a configuration of the geometric elements, $E$, that satisfies all these constraints. To achieve this, we define a loss function:
$L(E)=\sum_{i=1}^{m}c_{i}(E),$
where each constraint function $c_{i}(E)$ is continuous and differentiable with respect to the parameters of $E$ shown as Figure 3(d). When the loss reaches its minimum, i.e., $c_{i}(E^{*})=0$ for all $i$, where $E^{*}$ represents the desired configuration, all constraints are satisfied. It is essential that these constraint functions are differentiable to facilitate gradient-based optimization methods. It is worth noting that through our framework, users can easily define their own loss function according to the required constraints.
(3) Optimization of the constraints. Consequently, the geometric synthesizing task becomes a constrained minimization problem to determine the configuration $E^{*}$:
min $\quad L(E)$
subject to $\quad c_{i}\geq 0\text{ for }i=1\dots m.$
The optimization process converges when the loss reaches a threshold, at which point the configuration $E^{*}$ is considered the final geometry diagram, shown as Figure 3(e).
We visualized the process of the optimization of the GeoSDF in Fig. 2 with four IMO examples of the geometry diagrams. The leftmost diagrams represent the initial state, while the rightmost diagrams depict the final state after the optimization process. The initial state is set randomly. Through optimization, the SDF diagrams gradually evolve from a random configuration $E$ to $E^{*}$, ultimately satisfying the diagram’s constraints. The detailed synthesis processes are in the Appendix LABEL:app:Synthesize_Process.
During the optimization of the SDF, the loss function may converge to a local minimum, leading to a suboptimal configuration. This can cause the geometry shape $E^{*}$ to degrade into a degenerate form, such as a point or a line. In extreme cases, although the geometry may still satisfy most constraints (such as the perpendicularity of two lines), the resulting configuration will appear as a collapsed or overlapping set of elements, like a stack of lines or points, which is visually undesirable. The local minimum can be identified when the shape $E^{*}$ degenerates into a point or a line. To overcome this issue, we propose a crowd regularization term, which encourages a more spread-out configuration, preventing the elements from collapsing too close together. The regularization term is defined as
(2) $L_{crowd}(x)=\sum_{1\leq i<j\leq N}\left[\max\left(0,\tau-\|x_{i}-x_{j}\|\right)\right]^{2},$
where $\tau$ is the minimum acceptable distance between any pair of elements, and more details are in the Appendix LABEL:app:constraints.
@@
Table 1. Predefined common constraints and loss functions in GeoSDF. In our framework, GeoSDF, we consider these geometric constraints directly as the loss functions to quantify the deviation from the desired configuration. The mathematical details are in Appendix LABEL:app:constraints. Users could extend the extra loss function easily to satisfy other customized constraints.
Constraint Category Constraint Loss Function Description
@@
Incidence Constraints Equality $\text{Equal}(x_{1},x_{2})$ Two geometric elements $x_{1},x_{2}$ have the same value.
Less $\text{Less}(v_{1},v_{2})$ One numerical value is less than or equal to another.
@@
Metric Constraints Distance $\text{Equal}(E(x),v)$ A specific distance $v$ between two geometric elements $E_{1}(E_{2})$.
Angle $\text{Angle}(x_{1},x_{2},x_{3})$ The angle formed by three points $x_{1},x_{2},x_{3}$
Area $\text{Area}(x_{1},...,x_{n})$ A specific area for a polygon defined by a sequence of points.
Crowd Penalty $\text{Crowd}(x_{1},...,x_{n})$ Encourages geometric elements (typically points) to be spaced apart.
@@
Relative Position and Orientation Constraints Parallelism $\text{Parallel}(x_{1}x_{2},x_{3}x_{4})$ Two lines or segments are parallel to each other.
Perpendicularity $\text{Perpendicular}(x_{1}x_{2},x_{3}x_{4})$ Two line or segments are perpendicular to each other.
Order $\text{Order}(x_{1},...,x_{n})$ Counterclockwise order of points $x_{1},...,x_{n}$.
On $E(x)$ A specified point $x$ lies on a given segment or circle $E$.
@@
Inside $\text{Inside}(p,x_{1},...,x_{n})$ A point $p$ if it lies outside a convex polygon defined by other points.
@@
Topological Constraints Convexity $\text{Convex}(x_{1},...,x_{n})$ A polygon formed by a sequence of points $x_{1},...,x_{n}$ is convex.
(4) Rendering of diagrams. The explicit representation of the shape’s boundary is obtained by sampling the zero iso-surface of the implicit field, as depicted in Fig. 3 (f). This iso-surface is mathematically defined as the boundary of points $x$ where the signed distance function $F(x)=0$, thereby delineating the underlying planar geometry.
Figure 4. (a) The values sampled from SDF by $512\times 512$ pixel grid visualization. And visualizations of its zero iso-surface with different thresholds (b) 0.03, (c) 0.2, (d) 0.1, (e) 0.005.
4. Experiments

### v2

@@ Abstract
Plane Geometry Diagram Synthesis has been a crucial task in computer graphics, with applications ranging from educational tools to AI-driven mathematical reasoning. Traditionally, we rely on manual tools (e.g., Matplotlib and GeoGebra) to generate precise diagrams, but this usually requires huge, complicated calculations. Recently, researchers start to work on model-based methods (e.g., Stable Diffusion and GPT5) to automatically generate diagrams, saving operational cost but usually suffering from limited realism and insufficient accuracy. In this paper, we propose a novel framework GeoSDF, to automatically generate diagrams efficiently and accurately with Signed Distance Field (SDF). Specifically, we first represent geometric elements (e.g., points, segments, and circles) in the SDF, then construct a series of constraint functions to represent geometric relationships. Next, we optimize those constructed constraint functions to get an optimized field of both elements and constraints. Finally, by rendering the optimized field, we can obtain the synthesized diagram. In our GeoSDF, we define a symbolic language to represent geometric elements and constraints, and our synthesized geometry diagrams can be self-verified in the SDF, ensuring both mathematical accuracy and visual plausibility. In experiments, through both qualitative and quantitative analysis, GeoSDF synthesized both normal high-school level and IMO-level geometry diagrams. We achieve 88.67% synthesis accuracy by human evaluation in the IMO problem set. Furthermore, we obtain a very high accuracy of solving geometry problems (over 95% while the current SOTA accuracy is around 75%) by leveraging our self-verification property. All of these demonstrate the advantage of GeoSDF, paving the way for more sophisticated, accurate, and flexible generation of geometric diagrams for a wide array of applications.
1 Introduction
Figure 1: Comparison of geometric diagram synthesis methods for IMO 2022, Problem 4. (a) Rule-based systems, such as R-COT, rely on predefined templates and fail to produce diagrams when faced with non-predefined geometric constraints. (b) General model-based systems, including Gemini, DALL-E 3, and GPT-4o, often lack the precision for accurate mathematical synthesis, resulting in geometrically inconsistent figures. (c) In contrast, our proposed method accurately constructs the diagram by interpreting the fundamental geometric elements and constraints directly from the problem description.
@@ 1 Introduction
The study of geometry problems is fundamental across numerous disciplines, serving as a cornerstone for advancements in computer graphics Ye et al. (2020), computational geometry Dorst & Mann (2002), and even theoretical mathematics Trinh et al. (2024). The emergence of large language models (LLMs) has introduced new approaches for addressing Plane Geometry Problems (PGPs) Shi et al. (2024) by understanding both textual problem descriptions and their corresponding geometric diagrams Gao et al. (2025). Despite significant efforts to enhance the reasoning capabilities of LLMs, the performance of models remains limited by the scarcity of high-quality datasets containing annotated plane geometry diagrams Zhang et al. (2023b).
@@
In this paper, we propose a novel framework to automatically synthesize precise Geometry diagrams by leveraging Signed Distance Field (SDF), termed as GeoSDF. In the first step, problem descriptions are parsed into a structured symbolic representation consisting of elements (e.g., points, lines, circles) as well as the relationship among them, which are termed as constraints. We then represent each geometric element as a signed distance field (SDF), a continuous and differentiable function that compactly encodes its shape and position. Geometric constraints are expressed as differentiable functions over these fields, allowing us to formulate a global loss that enforces all constraints simultaneously. To prevent degenerate configurations such as overlapping elements, we additionally introduce a crowd regularization term that encourages well-spaced solutions. The resulting differentiable objective can be optimized using standard gradient-based methods. Finally, the optimized geometric elements are rendered into clear, human-readable diagrams, providing precise diagrams that faithfully reflect the mathematical relationships.
Previous works Ye et al. (2020) adopted a hard learning curve of domain-specific languages; in contrast, we leverage Large Language Models (LLMs) to translate natural language directly into a symbolic format. Based on simple use, our framework has the state-of-the-art performance on downstream tasks. We demonstrate the effectiveness on a range of challenging datasets, from the high-school level to an IMO-level benchmark. Our method achieves a problem-solving accuracy of 95.9%, a high-confidence result that confirms the reliability of our tool and surpasses the previous best by over 20 percentage points on the GeoQA dataset. And, our method enables the successful synthesis of diagrams for IMO-level problems where prior methods fall short. Moreover, our optimization process operates in batches, enabling the highly efficient, parallel synthesis of multiple diagrams.
The contributions of this paper are summarized as follows:
@@
  • We introduce GeoSDF, a novel and accurate framework for synthesizing plane geometry diagrams by optimizing SDF representation against symbolic mathematical constraints.
  • We design a more accessible user workflow where natural language is translated by an LLM into a symbolic representation, eliminating the need for specialized languages required by previous tools.
  • We achieve high efficiency through a batch-enabled optimization process, allowing for the simultaneous generation of multiple complex diagrams.
  • To demonstrate our method’s ability to generate high quantity and quality outputs, we augmented the existing data by synthesizing 32 diagrams with diversity for each problem description in FormalGeo7k dataset, finally forming a large size of dataset with 224k samples.
  • We achieve SOTA results on the GeoQA benchmark with 95.9% accuracy, demonstrating that high-quality diagram synthesis directly translates to superior geometric problem-solving capabilities.
2 Related Work
2.1 Plane Geometry Problem Solving
@@
Early approaches to plane geometry problem (PGP) solving relied on manually defined rules applied to small datasets Seo et al. (2015; 2014), resulting in poor generalization. Neural network-based models, such as NGS Chen et al. (2021) and DPE-NGS Chen et al. (2021), introduced visual question answering and specialized program generation, but still struggle with coarse-grained diagram understanding Ning et al. (2023). Symbolic reasoning methods, including Inter-GPS Lu et al. (2021) and FormalGeo Zhang et al. (2023b), use complex rule-based systems to interpret formal languages parsed from diagrams and text. However, their performance is constrained by limited datasets and parameter sizes, and they do not produce natural language solutions. Recent advances in multi-modal large language models (MLLMs), such as G-LLaVA Gao et al. (2025) and Geo-LLaVA Xu et al. (2024), fine-tune based models Liu et al. (2023) to generate natural language solutions. These approaches benefit from data augmentation techniques (e.g., Geo170K via GPT-3.5 Gao et al. (2025)), but often prioritize textual information over the diversity and complexity of geometry diagrams Zhuang et al. (2024). While methods like GeoX Xia et al. (2025) and R-CoT Deng et al. (2024) have made progress in diagram understanding, generating accurate, diverse, and controllable geometric diagrams remains a fundamental challenge for advancing PGP solving.
2.2 Plane Geometry Diagram Synthesis
@@
Plane geometry diagram synthesis is critical for evaluating problem-solving systems but is underexplored due to complexity Rombach et al. (2022). Current rule-based methods, including GeomVerse Kazemi et al. (2024) and MAVIS Zhang et al. (2025), stitch predefined shapes along edges. However, this limits them from generating diagrams with inscribed relationships (e.g., a triangle in a circle). Despite R-CoT’s Deng et al. (2024) advancements with inscribed elements, existing approaches remain constrained by predefined shapes and rules, preventing synthesis from textual requirements. Another approach is Penrose Ye et al. (2020), which synthesizes diagrams from a domain-specific language by solving a constrained numerical optimization problem. However, for problems in Euclidean geometry that demand high mathematical fidelity, its reliance on a general-purpose constraint solver can be less ideal. Our work differs fundamentally in its representation: GeoSDF leverages Signed Distance Fields to directly embed core geometric properties—like distance, angle, and relations—into the loss functions. This specialized approach results in a more stable and direct convergence to mathematically precise configurations.
3 Methodology
@@
The GeoSDF framework is developed for synthesizing plane geometry diagrams and consists of four steps, as illustrated in Fig. 2. Each step is described in detail in the following sections.
@@
Figure 2: The GeoSDF Pipeline with an example IMO diagram synthesis process. (a) Parsing text description in natural language to geometric Elements and Constraints in symbolic language. (b) The elements and constraints are represented using SDF representations. (c) The optimization process from the initialized field to satisfy all the specified constraints forms the optimized Fields. (d) The final optimized fields are rendered to Diagrams that visualize the geometric problem.
3.1 Parsing Natural Language into Symbolic Language
@@
Table 1: Predefined common constraints and loss functions in GeoSDF. In our framework, GeoSDF, we consider these geometric constraints directly as the loss functions to quantify the deviation from the desired configuration. The mathematical details are in Appendix A. Users could easily extend the extra loss function to satisfy other customized constraints.
Category Constraint Loss Function Description
@@
Incidence Equality $\text{Equal}(A,B)$ Two geometric elements $A,B$ have the same value.
Less $\text{Less}(A,B)$ One numerical value is less than or equal to another.
@@
Metric Distance $\text{Equal}(A(B),v)$ A specific distance $v$ between two geometric elements $A~\text{to}~B$.
Angle $\text{Angle}(A,B,C)$ The angle formed by three points $A,B,C$
Area $\text{Area}(A,...,E)$ A specific area for a polygon defined by a sequence of points.
@@
Spatial Parallelism $\text{Parallel}(AB,CD)$ Two lines or segments are parallel to each other.
Perpendicularity $\text{Perpendicular}(AB,CD)$ Two line or segments are perpendicular to each other.
Order $\text{Order}(A,...,E)$ Counterclockwise order of points A,,EA,...,E.
On $AB(C)$ A specified point $C$ lies on a given segment or circle $AB$.
@@
Inside $\text{Inside}(p,A,...,E)$ A point $p$ if it lies outside a convex polygon defined by other points $A$ to $E$.
@@
Topological Convexity $\text{Convex}(A,...,E)$ A polygon formed by a sequence of points A,,EA,...,E is convex.
The first parsing step of the GeoSDF framework converts geometric natural language statements into a structured symbolic language representation in an end-to-end manner. As illustrated in Fig. 2, the symbolic language consists of Elements, which denote fundamental geometric objects such as points, lines, or segments, and Constraints, which describe relationships or properties among these objects. Following previous works Berg et al. (2008), we consider four types of constraints: (1) incidence (e.g., a point lies on a circle), (2) metric (e.g., distance between two points equals a given value), (3) spatial (e.g., two lines are parallel), and (4) topological (e.g., convexity is preserved), concluded in Table 1.
For instance, consider a case where two line segments are defined by points A, B and C, D, respectively, and the two segments are parallel. After processing through this step, the structured symbolic representation encodes the elements as AB = SEGMENT(A, B) and CD = SEGMENT(C, D), while the constraint Parallel(AB, CD) captures the orientation relationship between the two segments. We fine-tune a QWen2.5-7B-Instruct model Qwen et al. (2025) on the training data in FormalGeo-7K Zhang et al. (2023b) to obtain the model for the parsing process.
3.2 Representation by SDF
We represent a geometric diagram, composed of the aforementioned elements and constraints, as a collection of signed distance fields (SDFs). An SDF is a continuous and differentiable scalar field that assigns to each point $(x,y)\in\mathbb{R}^{2}$ its shortest distance to a given geometric object. More details about SDF can be found in Appendix H.
Each element in the diagram is represented by its own SDF, which is fully determined by the element’s continuous parameters (e.g., coordinates for points, radii for circles, lengths for segments). This representation captures the element’s shape and position, providing a differentiable field that can later be optimized.
Constraints that encode geometric relationships between elements are expressed as differentiable functions of element parameters. They do not introduce new fields, rather, they specify conditions under which the combined element fields form a valid diagram. Each constraint thus acts as an optimization objective over the corresponding SDFs.
Finally, the complete diagram SDF is obtained by composing the individual element fields into a unified representation. As illustrated in Fig. 2, we begin by initializing a random field for each element. In the visualization, lighter colors indicate smaller SDF values, corresponding to points closer to the element. By sampling the field, we can observe the underlying structure of the initialized diagram. During optimization, the element parameters are adjusted to satisfy all constraints simultaneously, yielding a valid geometric configuration. We show a complete example in Appendix E.
3.3 Optimizing SDF Field
To clarify the optimization step, the diagram shows the complete set of $N$ geometric elements, $E=\{e_{1},e_{2},\dots,e_{N}\}$. Each element $e_{j}\in E$ is associated with its own set of differentiable constraints $C_{j}=\{c_{j,k}\mid k=1,2,\dots,M_{j}\}$, where $M_{j}$ is the number of constraints associated with element $e_{j}$, $k$ explicitly indexes the constraints for element $e_{j}$, so that $c_{j,k}$ denotes the $k$-th constraint in $C_{j}$.
The complete set of constraints across all elements is then $C=\bigcup_{j=1}^{N}C_{j}$. A valid geometric configuration is obtained when all constraints in $C$ are satisfied simultaneously. To enforce all geometric constraints, we have the loss function:
$L_{constraints}(E,C)=\sum_{j=1}^{N}\sum_{c_{j,k}\in C_{j}}c_{j,k}(E),$ (1)
where each $c_{j,k}$ is transformed into a non-negative differentiable function that evaluates to zero when the corresponding constraint is satisfied, ensuring the loss is bounded below by zero and suitable for gradient-based optimization.
To prevent elements from collapsing into overlapping or degenerate configurations, which could otherwise cause the optimization to fail, we introduce a crowd regularization term:
$L_{crowd}(E)=\sum_{1\leq i<j\leq N}\left[\max\left(0,\tau_{r}-\|x_{i}-x_{j}\|\right)\right]^{2},$ (2)
where $x_{i}$ and $x_{j}$ denote the position parameters of elements $e_{i}$ and $e_{j}$, and the hyper-parameter $\tau_{r}$ specifies a distance between any pair of elements.
The differentiable term can be optimized jointly with the SDF-based constraint losses. The total loss over all elements and constraints is then defined as:
$L_{total}=L_{constraints}(E,C)+\lambda L_{crowd}(E),$ (3)
where $\lambda$ balances the relative importance of the regularization. This differentiable loss is minimized using standard gradient-based optimization, yielding a configuration $E^{*}$ in which constraints are approximately satisfied and elements remain well-spaced. Further visualization of the optimization procedure is provided in Appendix B.
3.4 Boundary Extraction and Visualization
The optimized configuration $E^{*}$ specifies the final shapes of all geometric elements. Each element’s geometric parameters (e.g., center and radius of a circle) uniquely determine its boundary. Therefore, the boundaries can be directly computed from the elements representation $E^{*}$. Detailed boundary extraction procedures are provided in Appendix C.
4 Experiments